OpenStudy (anonymous):

Medal! Help please? A dealer has a lot that can hold 30 vehicles. In this lot, there are two available models, A and B. The dealer normally sells at least twice as many model A cars as model B cars. If the dealer makes a profit of $1300 on model A cars and $1700 on model B cars, how many of each car should the dealer have in the lot given that the total profit for the sale of both cars is $43,000?

OpenStudy (anonymous):

a. 20 model A 10 model B b. 10 model A 20 model B c. 5 model A 25 model B d. 16 model A 14 model B

OpenStudy (anonymous):

Oh, and I know that the answer is not B.

OpenStudy (anonymous):

I just went through and plugged all of the numbers into their given prices and got A. Model A: 1300(20) = 26000 Model B: 1700(10) = 17000 26000+17000= 43,000 but how would I create the equation/inequality to solve for this answer?

OpenStudy (tkhunny):

A + B <= 30 seems to be missing.

OpenStudy (tkhunny):

You might need A = 2B in there, but I'm a little worried about this one with the negative sign. What do we know of slack variables?

OpenStudy (anonymous):

Slack variables?

OpenStudy (tkhunny):

You'll meet them later, perhaps. For now, let's not worry about it. A = 2B results in A - 2B = 0. That should be a little unsettling as you may not have used "=" in this sort of a problem.

OpenStudy (anonymous):

Mkay, I following you.

OpenStudy (tkhunny):

Collect the stuff we've been looking at... A + B <= 30 A - 2B = 0 Maximize: 1300A + 1700B OR, since it's really the same problem, figure out how to get $43,000 out of the deal. a, b, c, and d all add up to 30, so that doesn't clear anything out. There is only one, a, where A = 2B. That may just be the solution.